# Constraining the Extragalactic Magnetic Field: Auger Data Meet UHECR Propagation Modeling

### v1

@@ 5 Summary and conclusions
We have here revisited a recent analysis by Auger which found a correlation between the UHECR arrival directions and the distribution of local SBGs or jetted AGNs (Abreu et al., 2022). Their analysis parameterized the deflections from EGMFs with a generic, source-distance independent “Fisher search radius”. The arrival directions of a fraction (“signal fraction”) of the full UHECR dataset, for particles with energies above an “energy threshold”, were found to correlate with the source catalog. The three parameters of their model (Fisher search radius, signal fraction, energy threshold) were then varied to maximize the test statistics.
We adopted the original Auger model as close as possible, but have replaced the parameter “Fisher search radius” with a more realistic EGMF model, taking into account the individual (distance-dependent) EGMF deflections of the individual sources. Subsequently, we calculated these deflections (and the UHECR energy losses) from detailed UHECR transport simulations, considering the expected UHECR spectrum and composition in the relevant energy range. Consequently, the corresponding new parameter we constrained is the turbulent EGMF strength $\tilde{B}$ in nG Mpc1/2, where the degeneracy with the coherence length is taken into account to first order. The other two parameters (signal fraction and energy threshold) were adopted with the same meaning as in Abreu et al. (2022).
As our main result, we have derived a 90% C.L. upper limit value on the EGMF. A constraint of 4.4 nG Mpc1/2 is obtained for the SBGs scenario, and 6.7 nG Mpc1/2 for the jetted AGN scenario. The corresponding upper limit on the turbulent GMF in the Galactic halo is dependent on the assumed halo size. Adopting a halo of size 30 kpc, an upper limit of $1.4~{}\mu$G kpc1/2 is obtained.
@@
The data sets generated and/or analyzed during this study are available from the corresponding author upon reasonable request. The UHECR propagation simulations were performed using the publicly available software package CRPropa 3 (Alves Batista et al., 2022). The availability of Auger’s software and data is described in Abreu et al. (2022) and is publicly accessible at https://doi.org/10.5281/zenodo.6504276.
@@ Acknowledgements
AAZ and AvV acknowledge support from Khalifa University’s internal FSU-2022-025 and RIG-2024-047 grants. AT and WW acknowledge support from DESY (Zeuthen, Germany), a member of the Helmholtz Association HGF. We wish to acknowledge the contribution of Khalifa University’s high-performance computing and research computing facilities to the results of this research. We would like to thank the Pierre Auger Collaboration for making their data and software used for Abreu et al. (2022) publicly available. We would like to thank the participants of the “Generation, evolution, and observations of cosmological magnetic fields” scientific program at the Bernoulli Center in April 2024, and the program “Towards a Comprehensive Model of the Galactic Magnetic Field” at Nordita in April 2023, which is partly supported by NordForsk the Royal Astronomical Society, for useful discussions.
@@ Appendix A Comparison with analytic estimates
A comparison between the simulation results and an analytical estimate of the angular deflection of high-energy particles due to turbulent extragalactic magnetic fields provides valuable insight into the accuracy and limitations of the analytical models used for predicting UHECR deflections. A simple analytical expression for the root-mean-square (rms) deflection angle $\theta_{\text{rms}}$ can be obtained assuming a set of consecutive small-angle deflections over its propagation distance $D$ through a turbulent field, the rms deflection angle $\theta_{\mathrm{rms}}$ can be approximated as (Harari et al., 2002; Bray & Scaife, 2018):

### v2

@@ 5 Summary and conclusions
We have here revisited a recent analysis by Auger which found a correlation between the UHECR arrival directions and the distribution of local SBGs or jetted AGNs (Abreu_2022). Their analysis parameterized the deflections from EGMFs with a generic, source-distance independent “Fisher search radius”. The arrival directions of a fraction (“signal fraction”) of the full UHECR dataset, for particles with energies above an “energy threshold”, were found to correlate with the source catalog. The three parameters of their model (Fisher search radius, signal fraction, energy threshold) were then varied to maximize the test statistics.
We adopted the original Auger model as close as possible, but have replaced the parameter “Fisher search radius” with a more realistic EGMF model, taking into account the individual (distance-dependent) EGMF deflections of the individual sources. Subsequently, we calculated these deflections (and the UHECR energy losses) from detailed UHECR transport simulations, considering the expected UHECR spectrum and composition in the relevant energy range. Consequently, the corresponding new parameter we constrained is the turbulent EGMF strength $\tilde{B}$ in nG Mpc1/2, where the degeneracy with the coherence length is taken into account to first order. The other two parameters (signal fraction and energy threshold) were adopted with the same meaning as in Abreu_2022.
[…]
@@
The data sets generated and/or analyzed during this study are available from the corresponding author upon reasonable request. The UHECR propagation simulations were performed using the publicly available software package CRPropa 3 (AlvesBatista_2022). The availability of Auger’s software and data is described in Abreu_2022 and is publicly accessible at https://doi.org/10.5281/zenodo.6504276.
@@ Acknowledgements
AAZ and AvV acknowledge support from Khalifa University’s internal FSU-2022-025 and RIG-2024-047 grants. AT and WW acknowledge support from DESY (Zeuthen, Germany), a member of the Helmholtz Association HGF. We wish to acknowledge the contribution of Khalifa University’s high-performance computing and research computing facilities to the results of this research. We would like to thank the Pierre Auger Collaboration for making their data and software used for Abreu_2022 publicly available. We would like to thank the participants of the “Generation, evolution, and observations of cosmological magnetic fields” scientific program at the Bernoulli Center in April 2024, and the program “Towards a Comprehensive Model of the Galactic Magnetic Field” at Nordita in April 2023, which is partly supported by NordForsk the Royal Astronomical Society, for useful discussions.
@@ Appendix A Comparison with analytic estimates
A comparison between the simulation results and an analytical estimate of the angular deflection of high-energy particles due to turbulent extragalactic magnetic fields provides valuable insight into the accuracy and limitations of the analytical models used for predicting UHECR deflections. A simple analytical expression for the root-mean-square (rms) deflection angle $\theta_{\text{rms}}$ can be obtained assuming a set of consecutive small-angle deflections over its propagation distance $D$ through a turbulent field, the rms deflection angle $\theta_{\mathrm{rms}}$ can be approximated as (Harari:2002dy; Bray_2018):